OpenStudy (anonymous):

Hi, we just started vectors in maths class and I don't understand a thing! Could someone please explain: Addition of 3 vectors coordinates Example: z = (-5, -8) u = (-2, -6) w = (-3, -6) z+u+w = ??? v = (6, 4) Also, the question says: prove algebraically that vectors v and w have the same direction but different magnitudes

OpenStudy (anonymous):

A vector is a direction. If all these vectors are with respect to the origin (0,0) then you simply have to plot the points, and connect the line. I used the v from your expression (it's obviously not int he same direction as your w, because it is wrong)

OpenStudy (anonymous):

to add vectors, you need to add teh components, (x with x, y with y) z + w + u = (-5,-8) + (-3, -6) + (-2, -6) = (-5 + -3 + -2, -8 + -6 + -6) = (-10, -20)=v

OpenStudy (anonymous):

A vector is usually written with coordinates, <x,y> how much it moves in the x direction, how much it moves in the y direction. w= <-3,-6> for every 3 it moves left, it moves 6 down. v= <-10,-20> for every 10 it moves left, it moves 20 down. Notice how w = 3* <-1,-2> and v = 10*<-1,-2> They have the same vectors! Only different multipliers or magnitude.

OpenStudy (anonymous):

Thanks so much could you also help me with this: w x ( v + z) w = -3, -2 v = 6, 4 z = -5, -8 In the answer book it says 5 but I didn't get that answer

OpenStudy (anonymous):

Oh yeah the x up there is meant to signify multiplication.

OpenStudy (anonymous):

when multiplying, you follow a similar concept. <x1,y1> * <x2, y2> = <x1 *x2, y1 * y2> you have to multiply the x with the x and the y with the y

OpenStudy (anonymous):

I did that but did not get the final answer "5"

OpenStudy (anonymous):

oh that's not an x, that's a cross. A cross product is the magnitude of the two vectors multiplied by sin ( theta) where theta is the angle between the two vectors.